OpenStudy (anonymous):

Given the rref(A)= U a 4x5 matrix and you are given column vectors of A: a1, a2 where rank(A) is 3 and a1, a2, and a4 are linearly independent. Is there a way to determine the remaining column vectors without using elementary row operations from U to a1, and a2?

OpenStudy (anonymous):

This is a purely conceptual question. I don't see fit to include the actual example matrices. So I was thinking performing the Row operations were very tedious and that in theory you should be able to write the remaining column vectors in some form of linear combination.

OpenStudy (anonymous):

I actually think that is the point of the question..

OpenStudy (anonymous):

@amistre64 : tell your friends, and their friends..

OpenStudy (amistre64):

i dont have any friends :(

OpenStudy (amistre64):

it seems like the remaining vectors would form a nul space

OpenStudy (anonymous):

Openstudy doesn't have a feature that lets you cash out your fans for friends? I mean you have 1,739 fans....you'd be rich with friends

OpenStudy (amistre64):

:)

OpenStudy (amistre64):

i forget what rank is in matrixes

OpenStudy (anonymous):

number of linearly independent columns

OpenStudy (anonymous):

rank(A)+dim(N(A))= n

OpenStudy (amistre64):

yeah, that sounds familiar. and the remaining columns define the nul space. in this case its either a plane of a line depending on if the remaining vectors are dependant of each other or not

OpenStudy (amistre64):

im not so sure theres a simple way to go about it other than row operations

OpenStudy (anonymous):

i only have column vectors a1, a2 and rref(A)=U. Since there are 3 independent vectors and I only have two of them is it possible to determine the other column vectors of A?

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

tedious, tedious

OpenStudy (amistre64):

matrixes are tedious ...

OpenStudy (anonymous):

Unfortunatly, I spent more time trying to find a way around the tediousness than actually just doing the operations.

OpenStudy (amistre64):

so:\[U=\begin{pmatrix} a_1&a_2&a_4&a_3&a_5\\ 1&0&0&c_1&c_2\\ 0&1&0&c_3&c_4\\ 0&0&1&c_5&c_6\\ 0&0&0&0&0 \end{pmatrix}\] \[\begin{matrix} a_1=&0+&-c_1 a_3& - c_2 a_5\\ a_2=&0+&-c_3 a_3& - c_4 a_5\\ a_3=&0+&1~a_3& +0~ a_5\\ a_4=&0+&-c_5 a_3& - c_6 a_5\\ a_5=&0+&+0~a_3& +1~a_5 \end{matrix}\]

OpenStudy (amistre64):

so, depending on what your 3 independant vectors are, and how it tediums to the rref, you can define the nulspace in terms of multiples of the dim vectors

OpenStudy (anonymous):

tediums?

OpenStudy (anonymous):

that could either be tedious or ties but not sure

OpenStudy (amistre64):

you say you are given 2 vectors, do you have to find the 3rd linearly independant one?

OpenStudy (anonymous):

A=original matrix, U=rref(A), and I have a1, a2 which are the column vectors of A. I need to find a3, a4, a5 for A

OpenStudy (amistre64):

give me the 2 vectors you got :)

OpenStudy (anonymous):

a1= (2,1,3,1)^T and a2=(-1,2,-3,1)^T

OpenStudy (anonymous):

U= [u1, u2, (2,3,0,0), u4, (-1,-2,5,0)] where u1,u2,u4 are lin. independent.

OpenStudy (anonymous):

obviously u1, u2,u4 contains pivotal 1s

OpenStudy (anonymous):

I'm honestly even having difficulty doing the row operations to find A

OpenStudy (amistre64):

\[A=\begin{pmatrix} a_1&a_2&a_4&a_3&a_5\\ 2&-1&k_1&n_1&n_2\\ 1&2&k_2&n_3&n_4\\ 3&-3&k_3&n_5&n_6\\ 1&1&k_4&n_7&n_8 \end{pmatrix}\] \[U=\begin{pmatrix} a_1&a_2&a_4&a_3&a_5\\ 1&0&0&c_1&c_2\\ 0&1&0&c_3&c_4\\ 0&0&1&c_5&c_6\\ 0&0&0&0&0 \end{pmatrix}\] \[\begin{pmatrix} 1&0&\frac{2k_1+k_2}{5}&\frac{2n_1+n_3}{5}&\frac{2n_2+n_4}{5}\\ 0&1&\frac{-k_1+2k_2}{5}&\frac{-n_1+2n_3}{5}&\frac{-n_2+2n_4}{5}\\ 0&0&1&\frac{-3n_1+6n_3-15n_1+10n_5}{-3k_1+6k_2-15k_1+10k_3}&\frac{-3n_2+6n_4-15n_2+10n_6}{-3k_1+6k_2-15k_1+10k_3}\\ 0&0&1&\frac{-3n_1+6n_3+5n_1-10n_7}{-3k_1+6k_2+5k_1-10k_4}&\frac{-3n_2+6n_4)+5n_2-10n_8}{-3k_1+6k_2+5k_1-10k_4} \end{pmatrix}\] \[2k_1+k_2=0=2k_2-k_1\] \[2k_1+k_2=4k_2-2k_1~;~k_2=0~:~k_1=0\]

OpenStudy (amistre64):

with any luck, i didnt mismath it along the way :)

OpenStudy (anonymous):

oh god, after all that work. I'd hate to just go: i don't understand what you just did.

OpenStudy (anonymous):

im thinking tho

OpenStudy (amistre64):

\[A=\begin{pmatrix} 2&-1&k_1\\ 1&2&k_2\\ 3&-3&k_3\\ 1&1&k_4 \end{pmatrix}\] \[\begin{pmatrix} 1&0&2k_4-k_2\\ 0&1&k_2-k_4\\ 0&0&\frac{k_1+3k_2-5k_4}{3}\\ 0&0&\frac{6k_2-9k_4+k_3}{6} \\ \end{pmatrix}\] \[k_1+3k_2-5k_4 = 3\] \[\begin{pmatrix} 1&0&2k_4-k_2\\ 0&1&k_2-k_4\\ 0&0&1\\ 0&0&\frac{6k_2-9k_4+k_3}{6} \\ \end{pmatrix}\] so, we have 4 equations and 4 unknowns \[k_1+3k_2-5k_4 = 3\\ 2k_4-k_2=0\\k_2-k_4=0\\6k_2-9k_4+k_3=0\] in order to turn this into\[\begin{pmatrix} 1&0&0\\ 0&1&0\\ 0&0&1\\ 0&0&0 \\ \end{pmatrix}\]

OpenStudy (amistre64):

im just doing row operations to reduce it to row echelon form

OpenStudy (amistre64):

from this i would say that k2=k4=0, k1=3, k3=0

OpenStudy (anonymous):

wow kudos! My opinion on this problem: stupid problem. If they were to give you (hypothetically) A(x)=b where x is some solution that equals some b. Couldn't you solve this problem way quicker?

OpenStudy (amistre64):

yes, with a guass jordon setup

OpenStudy (anonymous):

If I could thank you for your time - other than some virtual medal, I would.

OpenStudy (amistre64):

lol, im pretty much doing it for the practice, and the challenge :)

OpenStudy (amistre64):

and since a3 and a5 are dependant i would assume they form some sort of linear combination of a1, a2 and a4

OpenStudy (anonymous):

kudos for the exceptional attitude! hah

OpenStudy (amistre64):

... i broke the wolf trying to dbl chk myself :)

OpenStudy (anonymous):

ha

OpenStudy (anonymous):

Not to take up anymore time but I am thinking if you had some solution for the system x and the vector b that =Ax. You can essentially write the matrix A as (a1, a2, a4)x=b since a3, and a5 are dependent, is that right?

OpenStudy (amistre64):

yes, but that would not find a3 and a5

OpenStudy (anonymous):

oh, right

OpenStudy (anonymous):

wait so how would guass jordan setup go then? you'd have to write out the unknown vectors and just perform the operations and solve for the unknown after reducing?

OpenStudy (amistre64):

do you know if there is a specific solution? or is it such that there are many solutions and you need only find one?

OpenStudy (amistre64):

yes, that was one idea i had another idea is a grimm schmidy run

OpenStudy (anonymous):

grim schmidy? I know for sure I've never seen that.

OpenStudy (amistre64):

its a way to develop an orthogonal basis given 3 non orthoganal vectors

OpenStudy (anonymous):

oh, that must be further on in the course.

OpenStudy (amistre64):

also, a cross product will produce another vector that is linearly independant

OpenStudy (anonymous):

I've done calc 3 so it rings some bells

OpenStudy (anonymous):

completed/studied**

OpenStudy (amistre64):

another idea is to transpose the vectors into row vectors, add a vector <w,x,y,z>, or k subs if you like row reduce such that you get a last row of zeros ....

OpenStudy (anonymous):

Well, I must go chase Velocirapters out of my garden before they eat all the money on my money tree! Keep on keeping on, @amistre64 . I will be back with more tedious questions in due time.

OpenStudy (anonymous):

Cheers

OpenStudy (amistre64):

i dont think there is a unique results for this. assume that we know U, and lets pick some random stuff for the last 2 vectors\[ U=\begin{pmatrix} 1&0&0&1&1\\ 0&1&0&1&0\\ 0&0&1&1&1\\ 0&0&0&0&0 \end{pmatrix}\] we can then work "backwards" into A that contains the given vectors \[R_2+R_4=R_{new4} \begin{pmatrix} 1&0&0&1&1\\ 0&1&0&1&0\\ 0&0&1&1&1\\ 0&1&0&1&0 \end{pmatrix}\] \[-3R_2+R_3=R_{new3} \begin{pmatrix} 1&0&0&1&1\\ 0&1&0&1&0\\ 0&-3&1&-2&1\\ 0&1&0&1&0 \end{pmatrix}\] \[R_1+R_4=R_{new4} \begin{pmatrix} 1&0&0&1&1\\ 0&1&0&1&0\\ 0&-3&1&-2&1\\ 1&1&0&2&1 \end{pmatrix}\] \[3R_1+R_3=R_{new3} \begin{pmatrix} 1&0&0&1&1\\ 0&1&0&1&0\\ 3&-3&1&1&4\\ 1&1&0&2&1 \end{pmatrix}\] \[-\frac12R_2+R_1=R_{new1} \begin{pmatrix} 1&-1/2&0&1/2&1\\ 0&1&0&1&0\\ 3&-3&1&1&4\\ 1&1&0&2&1 \end{pmatrix}\] \[2R_1=R_{new1} \begin{pmatrix} 2&-1&0&1&2\\ 0&1&0&1&0\\ 3&-3&1&1&4\\ 1&1&0&2&1 \end{pmatrix}\] \[R_4+R_2=R_{new2} \begin{pmatrix} 2&-1&0&1&2\\ 1&2&0&3&1\\ 3&-3&1&1&4\\ 1&1&0&2&1 \end{pmatrix}\] therefore one possible setup for A is \[A= \begin{pmatrix} a_1&a_2&a_3&a_4&a_5\\ 2&-1&1&0&2\\ 1&2&0&0&1\\ 3&-3&1&1&4\\ 1&1&2&0&1 \end{pmatrix}\] the swap in columns then creates a slightly altered U form, but it still works http://www.wolframalpha.com/input/?i=rref%7B%7B2%2C-1%2C1%2C0%2C2%7D%2C%7B1%2C2%2C3%2C0%2C1%7D%2C%7B3%2C-3%2C1%2C1%2C4%7D%2C%7B1%2C1%2C2%2C0%2C1%7D%7D